# The WINTER Observatory: A One-Degree InGaAs Survey Camera to study the Transient Infrared Sky

### v1

@@ III.4.5 Linearity
To correct for these effects, we derive a per-pixel piecewise polynomial calibration. Each pixel is modeled with an 11th-order polynomial for the low-count region and a 5th-order polynomial for the high-count region, joined at a fixed crossover point of 25,000 counts with a 10,000-count overlap region. The fitting uses uniformly resampled exposure-time sequences of dome flats. During science processing, raw counts are passed through the appropriate segment of the piecewise model to yield an approximately linearized response, as seen in the dashed lines of Figure 9. The nonlinearity correction results in a median nonlinearity % across the WINTER sensors. The effect of these corrections on the science data is explored in V. Karambelkar and R. D. Stein et al., in prep.
@@
The nonlinearity correction coefficients, bad pixel masks, generation and implementation code are all released publicly through the winternlc package available on git.111https://github.com/winter-telescope/winternlc
@@ IV.1 Optical fly’s-eye design
WINTER’s optical design covers a 1 square degree field of view with 90% fill factor despite the sensors’ vacuum-sealed housing preventing close tiling (Figure 2). The novel fly’s-eye optical design splits the image between six identical optical channels to direct each section of the image onto its respective sensor with minimal inter-chip gaps between channels (Figure 10). To enable close packing of the optics, each lens is truncated symmetrically about one plane, matching the aspect ratio of the HD-format sensors. After passing through a waveband filter, the telescope’s focal plane is divided into six channels by a field-lens array made of six bonded lenses. A right-angle fold mirror then divides the image into two halves, placing three sensors on either end of the instrument to keep the center of gravity close to the telescope rotator (Hinrichsen:2020).
Furthermore, WINTER’s optical design prioritizes a 1″ pixel scale for the 15 m sensor pitch to match the median seeing at Palomar Observatory and other time-domain surveys such as the Palomar Transient Factory (PTF) and the Zwicky Transient Facility (ZTF). This necessitates a 2:1 demagnification of the F/6 telescope beam to an F/3 beam at the sensor. Each optical channel consists of ten spherical lenses. Lenses 1-5 collimate the telescope F/6 beam onto a pupil between lenses 5 and 6, and then lenses 6-10 reimage the pupil at an F/3 beam onto its respective InGaAs sensor.
All lens groups are bonded into groups of three tiled lens elements, with the exception of the field lens array (lens 1), which has six bonded lenses, and the lens 6-7 doublet, which retains independent adjustment to act as a compensator lens for fine focus optimization. Tolerancing studies simulating alignment procedures for the WINTER optical system for using Zemax’s Monte Carlo simulations found that maintaining strict fabrication tolerances for the lenses allows for loosening of mechanical alignment tolerances (Lourie:2020; Frostig:2020). This approach simplifies the optical alignment procedure, allowing for vertically tiled lens elements to be aligned together with the lens 6-7 bonded doublet allowing for independent adjustment for each optical channel. All WINTER lenses were purchased from Optimax Systems, who co-developed the precision truncation and alignment of the bonded, tiled lens elements, and all glasses were purchased from Ohara Corporation from the same melt batch to reduce variation between optical channels. To verify infrared performance, an outside company, M3, measured the index of refraction of a sample of each glass from 0.7 – 1.7 m. The infrared properties of these glasses had not yet been reported in the literature and closely matched predicted performance.
@@ IV.4 Optomechanics
The WINTER optomechanical design employs adhesive bonding for the lenses to reduce mass, design complexity, and stress on the optics. 3M Scotch-Weld 2216 A/B Gray epoxy was selected for all adhesive bonds based on extensive thermal and humidity lifetime tests simulating WINTER’s ten-year nominal lifetime (Hinrichsen:2020). The vertically-tiled lens elements are mounted in MIC-6 aluminum frames with flexure mounts made of Titanium 6Al-4V, providing single-axis lateral adjustment of the lens group (Figure 12). A custom 5-axis adjustable mount, made from Titanium 6Al-4V and Aluminum 7075-T6, holds the compensator lenses (6-7 doublet) with a stacked stage design where each stage adjusts one axis. The adjustment stages are accessible outside of the main instrument enclosure to provide flexibility for future focus modifications.
@@
To mitigate stray light, the optomechanical design also includes a focal plane mask and baffles around the tiled lens groups (Figure 12). The focal plane mask is placed in front of the field lens to prevent cross-talk between optical channels and reflections off of the bonded edges of the tiled lenses, which was identified as a major source of stray light in the non-sequential ray-tracing simulations outlined in (Frostig:2020). The focal plane mask is milled from MIC-6 cast aluminum and painted with Krylon Ultra-Flat Black, with the openings matching the footprints of the sensors on the telescope focal plane. Similar aluminum baffles painted with Aeroglaze Z306 flat black are mounted to the tiled lens group frames, with openings oversized relative to the image footprint. Prototype focal plane masks made with selective laser sintering (SLS) 3D printers were produced as well. The 3D printed versions captured the complex geometry of the baffle but did not achieve the knife-edge sharpness of the machined masks. While these were not optically characterized, with further optimization this approach may be viable for future efforts.
@@ IV.5 Alignment, integration, and test
WINTER’s custom truncated lenses require special alignment, integration, and test (AIT) procedures. In this section, we provide a concise overview of the AIT procedure, while a comprehensive analysis of the instrument AIT process and outcomes will appear in an upcoming publication. First, the as-manufactured lens thicknesses were measured with a TriOpics OptiSurf interferometer, prompting a re-optimization of the lens spacings based on the tolerancing procedures outlined in Lourie:2020; Frostig:2020. Next, the fold mirror was installed in the center of the instrument with a Hexagon coordinate measuring machines (CMM) arm used to verify the placement and tilt of the optic to a 25 m accuracy. Small mirrors were attached to the sensor mounts on either end of the instrument and aligned with a Zygo Verifire wavefront interferometer to pre-align the tip, tilt, and focus position of the sensors before delivery from the vendor. The tiled lens group were then aligned with the CMM arm, with the aluminum frames resting against reference pins and shimmed to compensate for tip and tilt along the optical axis. The individually-mounted compensator lenses were then aligned with their five-axis mounts and verified with the CMM arm. For the final alignment, a fiber-fed pinhole mask manufactured by FiberTech Optica provided an array of five F/3 sources per sensor, allowing for fine optical three-axis positioning of the sensors and verifying the required instrument image quality. Due to time constraints and the initial alignment meeting design requirements, a final optical alignment with compensator lenses was not conducted.
@@ IV.6 Camera housing/mechanics
A precision optical bench serves as the reference plane for aligning and mounting all optomechanical components, including the sensor mounting frames (Figure 13). The optical bench is made of a cast aluminum plate with flatness and parallelism tolerances of 0.025 mm and placement tolerance for mounting holes and alignment pins of 0.18 mm. It is painted with Aeroglaze Z306 flat black paint between non-critical surfaces to reduce stray light. Black-painted panels, resting on an aluminum frame and sealed with rubber gaskets to prevent dust contamination, form the remaining five sides of the enclosure. A focal plane mask located in front of the field lens array mitigates stray light due to cross-talk between optical channels (Figure 12). Similar baffles around the tiled lens arrays act as stray-light reduction and a safety retaining system in case of adhesive bond failure (Figure 12). To further prevent stray light near the image planes, Acktar Advanced Coatings’ Metal Velvet Black Foil was applied to the sensor mounting frames and parts of the filter tray. A custom telescope adapter made from carbon fiber reinforced plastic manufactured by CarbonVision mounts the instrument to the telescope rotator.
@@ V Observatory
WINTER operates at Palomar Observatory in Southern California, sharing a site with the 200-inch Hale Telescope, the Zwicky Transient Facility (ZTF) (ztf2019) and Palomar Gattini IR (De:2019xhw). The site is at an altitude of 1700 meters and is run by Caltech Optical Observatories. WINTER takes advantage of existing infrastructure at the site to run robotic telescopes, including data downlink, weather monitoring, and facilities operations. The weather and observing conditions are continuously monitored by telescope operators and support astronomers at the 200-inch Hale Telescope, who enable and lockout observing throughout the night accordingly. This continuous staffing support also allows for prompt hands-on response to operational anomalies and equipment failures that may occur over years of operation.
The instrument operates in a preexisting five-meter dome with a new, adjacent climate-controlled shed housing the WINTER control computers and electronics, power supplies, and chiller (Figure 15). The dome was retrofit for robotic operations, including a robotic control and telemetry system similar to those developed for ZTF and PGIR, a new weather station, motor drives, and a custom telescope pier. The dome’s robotic control system allows for extensive autonomous safety protocols to protect both workers and equipment, as outlined in Frostig:2020.
@@ V.1 Telescope
WINTER employs a 1-meter commercial off-the-shelf (COTS) corrected Dall-Kirkham (CDK) telescope made by PlaneWave Instruments with modifications for the project (Figure 15). The PW1000 1-Meter Observatory System features an instrument rotator mount to correct for sky rotation and fast direct-drive altitude and azimuth motors with a maximum velocity of 15 degrees per second with a settle time between slews of one second. The telescope optics include a 1-meter primary mirror, a spherical secondary mirror, a flat tertiary mirror with an integrated rotator to select between the two Nasmyth ports, and a set of three corrector lenses at each of the ports. All of the telescope optics are made of fused silica to increase thermal stability. Customizations for the WINTER project include developing a focusing system at the secondary mirror instead of the default focus mechanism at the instrument rotator mount, and applying near-infrared-optimized anti-reflection coatings on the corrector lenses for one Nasmyth port. The opposing port retains optical-wavelength-optimized coatings and houses a companion instrument.
@@ V.2 SUMMER and SPRING
The telescope and the companion instrument, SUMMER (Studying the Universe with Multi-Messenger and Exoplanet Research), were installed in June of 2021, ahead of WINTER’s completion in 2023, enabling development of the robotic control code, scheduler, and data reduction pipeline (see sections VI.1, VI.2, VI.3) and performing optical follow up of science targets (e.g., Christos:2023). The instrument consists of a Raptor Photonics CCD with e2v 42-40 architecture, which includes a 5-stage TEC and liquid cooling loop, a 65 mm optical shutter from Vincent Associates, and a Finger Lakes Instrumentation seven-position filter wheel which includes u, g, r, and i SDSS passband filters from Asahi Spectra. SUMMER images cover a 0.29-degree-squared field of view.
Since commissioning the telescope, the second port on the 1-meter telescope also served as a testbed for new sensors. There have been multiple observing runs with optical Complementary Metal-Oxide-Semiconductor (CMOS) sensors, including the Hamamatsu Photonics qCMOS camera and the large-format Teledyne COSMOS camera (see results in Layden:2025). Currently, it is home to the SPRING (Super PalomaR INGaas) camera, a Princeton Infrared Technologies 1280SciCam InGaAs camera with 1280 x 1024 12.5m pixels used for deep infrared observations of targets that fit within the smaller 8.8 x 7.0 arcminute field of view.
@@ VI.1 Robotic control software
The WINTER telescope’s operations are managed by the WINTER Supervisory Program (WSP), a highly parallelized Python 3.9-based control system leveraging PyQt5222https://riverbankcomputing.com/software/pyqt/ for interfacing with the Qt application framework333https://www.qt.io/development/qt-framework, enabling sophisticated control and scheduling capabilities (wsp_software_spie) 444https://github.com/winter-telescope/observatory/. The WSP consists of more than a dozen individual subsystem daemons which communicate with each other over TCP/IP sockets using the Pyro5 python library555pyro5.readthedocs.io, each of which are multi-threaded to allow concurrent commanding, communication, and housekeeping monitoring. The system is launched and managed as a Linux system daemon on the observatory’s control PC, with individual subsystem daemons run on separate PCs. WSP generates nightly schedules and manages observation tasks. This system, drawing upon methodologies from the MINERVA (MINERVA_Swift_2015), Robo-AO (robo-ao_Riddle_2012), GOTO (goto_dyer) and SuperBIT (superbit_romualdez) projects, facilitates automated decision-making for observatory operations, such as dome opening and observation initiation, while allowing manual override by operators in case of adverse conditions. The WSP orchestrates a variety of processes including scheduling, target of opportunity (ToO) event handling, command execution via a priority queue system, and comprehensive logging of telescope states and observations through real time telemetry graphing and slack notifications.
@@ VI.2 Scheduler and ToOs
To autonomously balance multiple observing programs with a range of cadences, filters, and priorities, WINTER employs a custom scheduling software based on the ZTF scheduler (ztf_scheduler_Bellm_2019). It creates nightly schedules based on WINTER’s surveys and long-term targets, which can be interrupted by Target of Opportunity (ToO) requests at any point in the night. A flexible API, winterapi666https://github.com/winter-telescope/winterapi allows authorized users to submit ToO requests directly to the queue, download data associated with their observing programs, and enables high-level tracking by the team to ensure adequate use of observing time. This framework allows observers to manage their own programs, and enables automated follow-up triggering of transient candidates without the need for an observer in the loop. The scheduler prioritizes programs based on their observational history such that any observations interrupted by weather or ToOs are rescheduled for subsequent nights. The WINTER scheduler divides the accessible northern sky ( declination) into a fixed pointing grid of fields matching the WINTER field of view. Each potential observation is prioritized by a volumetric weighing metric and then scheduled with one of three modes: Gurobi-optimized observing, greedy observing, and queue observing. The volumetric metric maximizes the limiting volume of a requested field at a given altitude, which is based on detailed simulations of the limiting magnitude including seeing, sky brightness, and instrument noise. For nightly surveys, Gurobi-optimized observing utilizes the Gurobi linear optimizer to solve a traveling salesman problem, balancing various scientific surveys with unique cadences while minimizing overhead time from telescope slews. Alternatively, greedy observing continually selects the best field based on current volumetric weighting and is used to fill in reference sky images between scheduled science surveys and ToOs. Finally, for any ToO requests, queue observing sequentially steps through a predefined program. Figure 16 shows WINTER observations from the first two years of observing, balancing various surveys as discussed in Section VII.
@@
For robotic follow up of gravitational-wave alerts, WINTER uses the Growth ToO Marshall (GROWTH) to ingest alerts distributed via the LIGO-Virgo GCN system and produce optimized observing schedules with the gwemopt package (Coughlin:2018lta; Coughlin:2019qkn; Ghosh:2015sxp). Frostig:2022 outlines extensive simulations of WINTER observing with the gwemopt package, leading to a tailored gravitational-wave follow up strategy based on the reported distance and localization area of each gravitational-wave event. WINTER only follows up events with at least a 10 chance of kilonova detection based on simulations with two kilonova model grids. For each qualifying event, WINTER dynamically adjusts the image exposure time to optimize the search strategy, dedicating up to seven nights searching for each event and repeatedly observing kilonova candidates until they fade.

### v2

@@ III.4.5 Linearity
To correct for these effects, we derive a per-pixel piecewise polynomial calibration. Each pixel is modeled with an 11th-order polynomial for the low-count region and a 5th-order polynomial for the high-count region, joined at a fixed crossover point of 25,000 counts with a 10,000-count overlap region. The fitting uses uniformly sampled exposure-time sequences of dome flats. During science processing, raw counts are passed through the appropriate segment of the piecewise model to yield an approximately linearized response, as seen in the dashed lines of Figure 9. The nonlinearity correction results in a median nonlinearity % across the WINTER sensors. The effect of these corrections on the science data is explored in V. Karambelkar and R. D. Stein et al., in prep.
@@
The nonlinearity correction coefficients, bad pixel masks, generation, and implementation code are all released publicly through the winternlc package available on git.111https://github.com/winter-telescope/winternlc
@@ IV.1 Optical fly’s-eye design
WINTER’s optical design covers a 1 deg2 field of view with 90% fill factor despite the sensors’ vacuum-sealed housing preventing close tiling (Figure 2). The novel fly’s-eye optical design splits the image between six identical optical channels to direct each section of the image onto its respective sensor with minimal inter-chip gaps between channels (Figure 10). To enable close packing of the optics, each lens is truncated symmetrically about one plane, matching the aspect ratio of the HD-format sensors. After passing through a waveband filter, the telescope’s focal plane is divided into six channels by a field-lens array made of six bonded lenses. A right-angle fold mirror then divides the image into two halves, placing three sensors on either end of the instrument to keep the center of gravity close to the telescope rotator (Hinrichsen et al., 2020).
Furthermore, WINTER’s optical design prioritizes a 1″ pixel scale for the 15 m sensor pitch to match the median seeing at Palomar Observatory and other time-domain surveys such as PTF and ZTF. This necessitates a 2:1 demagnification of the F/6 telescope beam to an F/3 beam at the sensor. Each optical channel consists of ten spherical lenses. Lenses 1-5 collimate the telescope F/6 beam onto a pupil between lenses 5 and 6, and then lenses 6-10 reimage the pupil at an F/3 beam onto its respective InGaAs sensor.
All lens groups are bonded into groups of three tiled lens elements, with the exception of the field lens array (lens 1), which has six bonded lenses, and the lens 6-7 doublet, which retains independent adjustment to act as a compensator lens for fine focus optimization. Tolerancing studies simulating alignment procedures for the WINTER optical system for using Zemax’s Monte Carlo simulations found that maintaining strict fabrication tolerances for the lenses allows for loosening of mechanical alignment tolerances (Lourie et al., 2020; Frostig et al., 2020). This approach simplifies the optical alignment procedure, allowing for vertically tiled lens elements to be aligned together with the lens 6-7 bonded doublet allowing for independent adjustment for each optical channel. All WINTER lenses were purchased from Optimax Systems, who co-developed the precision truncation and alignment of the bonded, tiled lens elements, and all glasses were purchased from Ohara Corporation from the same melt batch to reduce variation between optical channels. To verify infrared performance, an outside company, M3, measured the index of refraction of a sample of each glass from 0.7 – 1.7 m. The infrared properties of these glasses had not yet been reported in the literature and closely matched predicted performance.
@@ IV.4 Optomechanics
The WINTER optomechanical design employs adhesive bonding for the lenses to reduce mass, design complexity, and stress on the optics. 3M Scotch-Weld 2216 A/B Gray epoxy was selected for all adhesive bonds based on extensive thermal and humidity lifetime tests simulating WINTER’s ten-year nominal lifetime (Hinrichsen et al., 2020). The vertically-tiled lens elements are mounted in MIC-6 aluminum frames with flexure mounts made of Titanium 6Al-4V, providing single-axis lateral adjustment of the lens group (Figure 12). A custom 5-axis adjustable mount, made from Titanium 6Al-4V and Aluminum 7075-T6, holds the compensator lenses (6-7 doublet) with a stacked stage design where each stage adjusts one axis. The adjustment stages are accessible outside of the main instrument enclosure to provide flexibility for future focus modifications.
@@
To mitigate stray light, the optomechanical design also includes a focal plane mask and baffles around the tiled lens groups (Figure 12). The focal plane mask is placed in front of the field lens to prevent cross-talk between optical channels and reflections off of the bonded edges of the tiled lenses, which was identified as a major source of stray light in the non-sequential ray-tracing simulations outlined in Frostig et al. (2020). The focal plane mask is milled from MIC-6 cast aluminum and painted with Krylon Ultra-Flat Black, with the openings matching the footprints of the sensors on the telescope focal plane. Similar aluminum baffles painted with Aeroglaze Z306 flat black are mounted to the tiled lens group frames, with openings oversized relative to the image footprint. Prototype focal plane masks made with selective laser sintering (SLS) 3D printers were produced as well. The 3D printed versions captured the complex geometry of the baffle but did not achieve the knife-edge sharpness of the machined masks. While these were not optically characterized, with further optimization this approach may be viable for future efforts.
@@ IV.5 Alignment, integration, and test
WINTER’s custom truncated lenses require special alignment, integration, and test (AIT) procedures. In this section, we provide a concise overview of the AIT procedure, while a comprehensive analysis of the instrument AIT process and outcomes will appear in an upcoming publication. First, the as-manufactured lens thicknesses were measured with a TriOpics OptiSurf interferometer, prompting a re-optimization of the lens spacings based on the tolerancing procedures outlined in Lourie et al. (2020); Frostig et al. (2020). Next, the fold mirror was installed in the center of the instrument with a Hexagon coordinate measuring machines (CMM) arm used to verify the placement and tilt of the optic to a 25 m accuracy. Small mirrors were attached to the sensor mounts on either end of the instrument and aligned with a Zygo Verifire wavefront interferometer to pre-align the tip, tilt, and focus position of the sensors before delivery from the vendor. The tiled lens group were then aligned with the CMM arm, with the aluminum frames resting against reference pins and shimmed to compensate for tip and tilt along the optical axis. The individually-mounted compensator lenses were then aligned with their five-axis mounts and verified with the CMM arm. For the final alignment, a fiber-fed pinhole mask manufactured by FiberTech Optica provided an array of five F/3 sources per sensor, allowing for fine optical three-axis positioning of the sensors and verifying the required instrument image quality. Due to time constraints and the initial alignment meeting design requirements, a final optical alignment with compensator lenses was not conducted.
@@ IV.6 Camera housing/mechanics
A precision optical bench serves as the reference plane for aligning and mounting all optomechanical components, including the sensor mounting frames (Figure 13). The optical bench is made of a cast aluminum plate with flatness and parallelism tolerances of 0.025 mm and placement tolerance for mounting holes and alignment pins of 0.18 mm. It is painted with Aeroglaze Z306 flat black paint between non-critical surfaces to reduce stray light. Black-painted panels, resting on an aluminum frame and sealed with rubber gaskets to prevent dust contamination, form the remaining five sides of the enclosure. A focal plane mask located in front of the field lens array mitigates stray light due to cross-talk between optical channels (Figure 12). Similar baffles around the tiled lens arrays act as stray-light reduction and a safety retaining system in case of adhesive bond failure (Figure 12). To further prevent stray light near the image planes, Acktar Advanced Coatings’ Metal Velvet Black Foil was applied to the sensor mounting frames and parts of the filter tray. A custom carbon fiber–reinforced plastic adapter, manufactured by CarbonVision, mounts the instrument to the telescope rotator.
@@ V Observatory
WINTER operates at Palomar Observatory in Southern California, sharing a site with the 200-inch Hale Telescope, the Zwicky Transient Facility (ZTF) (Bellm et al., 2019b) and Palomar Gattini IR (De et al., 2020). The site is at an altitude of 1700 meters and is run by Caltech Optical Observatories. WINTER takes advantage of existing infrastructure at the site to run robotic telescopes, including data downlink, weather monitoring, and facilities operations. The weather and observing conditions are continuously monitored by telescope operators and support astronomers at the 200-inch Hale Telescope, who enable and lockout observing throughout the night accordingly. This continuous staffing support also allows for prompt hands-on response to operational anomalies and equipment failures that may occur over years of operation.
The instrument operates in a preexisting five-meter dome with a new, adjacent climate-controlled shed housing the WINTER control computers and electronics, power supplies, and chiller (Figure 15). The dome was retrofit for robotic operations, including a robotic control and telemetry system similar to those developed for ZTF and PGIR, a new weather station, motor drives, and a custom telescope pier. The dome’s robotic control system allows for extensive autonomous safety protocols to protect both workers and equipment, as outlined in Frostig et al. (2020).
@@ V.1 Telescope
WINTER employs a 1-meter commercial off-the-shelf (COTS) corrected Dall-Kirkham (CDK) telescope made by PlaneWave Instruments with modifications for the project (Figure 15). The PW1000 1-Meter Observatory System features an instrument rotator mount to correct for sky rotation and fast direct-drive altitude and azimuth motors with a maximum velocity of 15 degrees per second with a settle time between slews of one second. The telescope optics include a 1-meter primary mirror, a spherical secondary mirror, a flat tertiary mirror with an integrated rotator to select between the two Nasmyth ports, and a set of three corrector lenses at each of the ports. All of the telescope optics are made of fused silica to increase thermal stability. Customizations for the WINTER project include developing a focusing system at the secondary mirror instead of the default focus mechanism at the instrument rotator mount, and applying near-IR-optimized anti-reflection coatings on the corrector lenses for one Nasmyth port. The opposing port retains optical-wavelength-optimized coatings and houses a companion instrument.
@@ V.2 SUMMER and SPRING
The telescope and a companion instrument, SUMMER (Studying the Universe with Multi-Messenger and Exoplanet Research), were installed in June of 2021, ahead of WINTER’s completion in 2023, enabling development of the robotic control code, scheduler, and data reduction pipeline (see sections VI.1, VI.2, VI.3) and performing optical follow up of science targets (e.g., Panagiotou et al. (2023)). The instrument consisted of a Raptor Photonics CCD with e2v 42-40 architecture, which includes a 5-stage TEC and liquid cooling loop, a 65 mm optical shutter from Vincent Associates, and a Finger Lakes Instrumentation seven-position filter wheel which includes u, g, r, and i SDSS passband filters from Asahi Spectra. SUMMER images cover a 0.29-degree-squared field of view.
Since commissioning the telescope, the second port on the 1-meter telescope has also served as a testbed for new sensors. There have been multiple observing runs with optical Complementary Metal-Oxide-Semiconductor (CMOS) sensors, including the Hamamatsu Photonics qCMOS camera and the large-format Teledyne COSMOS camera (see results in Layden et al. 2025). Currently, it is home to the SPRING (Super PalomaR INGaas) camera, a Princeton Infrared Technologies 1280SciCam InGaAs camera with 1280 x 1024 12.5 m pixels used for deep infrared observations of targets that fit within the smaller 8.8 x 7.0 arcminute field of view.
@@ VI.1 Robotic control software
The WINTER telescope’s operations are managed by the WINTER Supervisory Program (wsp), a highly parallelized Python 3.9-based control system leveraging PyQt5222https://riverbankcomputing.com/software/pyqt/ for interfacing with the Qt application framework333https://www.qt.io/development/qt-framework, enabling sophisticated control and scheduling capabilities (Lourie et al., 2024) 444https://github.com/winter-telescope/observatory/. The WSP consists of more than a dozen individual subsystem daemons which communicate with each other over TCP/IP sockets using the Pyro5 python library555pyro5.readthedocs.io, each of which are multi-threaded to allow concurrent commanding, communication, and housekeeping monitoring. The system is launched and managed as a Linux system daemon on the observatory’s control PC, with individual subsystem daemons run on separate PCs. wsp generates nightly schedules and manages observation tasks. This system, drawing upon methodologies from the MINERVA (Swift et al., 2015), Robo-AO (Riddle et al., 2012), GOTO (Dyer et al., 2018) and SuperBIT (Romualdez et al., 2018) projects, facilitates automated decision-making for observatory operations, such as dome opening and observation initiation, while allowing manual override by operators in case of adverse conditions. wsp orchestrates a variety of processes including scheduling, Target of Opportunity (ToO) event handling, command execution via a priority queue system, and comprehensive logging of telescope states and observations through real time telemetry graphing and slack notifications.
@@ VI.2 Scheduler and ToOs
To autonomously balance multiple observing programs with a range of cadences, filters, and priorities, WINTER employs a custom scheduling software based on the ZTF scheduler (Bellm et al., 2019a). It creates nightly schedules based on WINTER’s surveys and long-term targets, which can be interrupted by ToO requests at any point in the night. A flexible API, winterapi666https://github.com/winter-telescope/winterapi allows authorized users to submit ToO requests directly to the queue, download data associated with their observing programs, and enables high-level tracking by the team to monitor observing time. This framework allows observers to manage their own programs, and enables automated follow-up triggering of transient candidates without the need for an observer in the loop. The scheduler prioritizes programs based on their observational history such that any observations interrupted by weather or ToOs are rescheduled for subsequent nights. The WINTER scheduler divides the accessible northern sky ( declination) into a fixed pointing grid of fields matching the WINTER field of view. Each potential observation is prioritized by a volumetric weighing metric and then scheduled with one of three modes: Gurobi-optimized observing, greedy observing, and queue observing. The volumetric metric maximizes the limiting volume of a requested field at a given altitude, which is based on detailed simulations of the limiting magnitude including seeing, sky brightness, and instrument noise. For nightly surveys, Gurobi-optimized observing utilizes the Gurobi linear optimizer to solve a traveling salesman problem, balancing various scientific surveys with unique cadences while minimizing overhead time from telescope slews. Alternatively, greedy observing continually selects the best field based on current volumetric weighting and is used to fill in reference sky images between scheduled science surveys and ToOs. Finally, for any ToO requests, queue observing sequentially steps through a predefined program. Figure 16 shows WINTER observations from the first two years of observing, balancing various surveys as discussed in Section VII.
@@
For robotic follow up of gravitational-wave alerts, WINTER uses the Growth ToO Marshall (Kasliwal et al., 2019) to ingest alerts distributed via the LIGO-Virgo GCN system and produce optimized observing schedules with the gwemopt package (Coughlin and others, 2018, 2019; Ghosh et al., 2016). Frostig et al. (2022a) outlines extensive simulations of WINTER observing with the gwemopt package, leading to a tailored gravitational-wave follow up strategy based on the reported distance and localization area of each gravitational-wave event. WINTER only follows up events with at least a 10 chance of kilonova detection based on simulations with two kilonova model grids. For each qualifying event, WINTER dynamically adjusts the image exposure time to optimize the search strategy, dedicating up to seven nights searching for each event and repeatedly observing kilonova candidates until they fade.